### v2

@@ 3.3 Light curve
The OSIRIS instrument was also used to study the rotational properties of the object. On September 28, 2024, a series of continuous 30 s exposures were taken during one hour using the Sloan- filter and the standard CCD configuration ( binning, plate scale of 0.254 ″ pixel-1, readout time of 21 s). At the time of observation, the object had an apparent magnitude and a total apparent angular rate of 1.2 ″ min-1. Standard bias and flat-field corrections were applied to reduce the images.
Relative aperture photometry was performed on the images using three field stars as reference stars. Different aperture radii were tested, including both fixed and variable apertures with aperture corrections. The best result, obtained using a fixed aperture radius of about FWHM, is shown in Fig. 4.
The signal-to-noise ratio (SNR) of the light curve is not sufficient to unambiguously determine a rotation period. However, a variability with an amplitude of at least 0.5 mag is observed during the one-hour observing block, which is consistent with the uncertainty in in Table 1. Therefore, we consider that the object may have a rotation period of around one hour or less. This would be similar to that of Kamo’oalewa, which has an estimated rotation period of min (Sharkey et al. 2021). For an object with an absolute magnitude of 27.4, the possibility of a higher rotational frequency or even tumbling motion cannot be excluded. However, observations with higher time resolution and better SNR are required to confirm any of such behaviors.
@@ 4 Discussion
Our spectroscopic results on the surface composition of 2024 PT5 are not conclusive due to the lack of a near-infrared reflectance spectra, but they are suggestive of a Lunar origin. Within the context of the Mission Accessible Near-Earth Object Survey (MANOS, Kareta et al. 2022), Kareta et al. used the Lowell Discovery Telescope (LDT) and the NASA Infrared Telescope Facility (IRTF) to obtain the visible and near-infrared spectrum of 2024 PT5 to find that it is well matched by samples of the Moon.101010https://iawn.net/documents/20240917_19th_virtual/IAWN_2024_Kareta.pdf As pointed out above, the available data indicate that (469219) 2016 HO3 Kamo’oalewa, 2022 NX1, and 2024 PT5 may all have a common origin as Lunar ejecta. The topic of the dynamics of Lunar ejecta in the neighborhood of the Earth–Moon system was first studied by Warren (1994) and Gladman et al. (1995) arriving at the conclusion that these materials can be scattered out of their source region relatively quickly, in an effective timescale of 105 yr. This subject is now receiving renewed attention (Castro-Cisneros et al. 2023; Jiao et al. 2024) after the confirmation of Kamo’oalewa as Lunar ejecta, but it is also being linked to the origin of the NEOs involved in temporary capture events (Jedicke et al. 2024). Nevertheless, the theoretical expectations drawn from the literature indicate that Kamo’oalewa, 2022 NX1, and 2024 PT5 may have all emerged out of lunar impact craters formed during the last Myr or so.
@@ 5 Summary and conclusions
In this Letter, we presented a detailed analysis of a mini-moon engagement of the temporarily captured flyby type experienced by Arjuna asteroid 2024 PT5 in 2024, from September 29 to November 25. The study of the short-term orbital evolution of 2024 PT5 was carried out using direct -body simulations. We derived its spectral class from a visible reflectance spectrum obtained with the OSIRIS camera spectrograph at the 10.4 m Gran Telescopio Canarias. Data from the same instrument were used to perform a preliminary assessment of its rotational state. Astrometry was obtained using the Two-meter Twin Telescope (see Appendix D). Our conclusions can be summarized as follows.
@@
We confirm that 2024 PT5 is a natural object with a visible spectrum consistent with that of an Sv-type asteroid, although it could also be classified as a Lunar mare breccia.
@@
We confirm that 2024 PT5 experienced a temporarily captured flyby between September 29 and November 25, 2024.
@@
The orbital evolutions of 2022 NX1 and 2024 PT5 are very similar as well as the circumstances of their temporarily captured flyby episodes.
Both ESA and NASA now emphasize a low-cost approach to NEO missions for small body science and planetary defense, including reuse and repurposing of already active missions (Freeman et al. 2024). The discovery and characterization of accessible objects like 2022 NX1 and 2024 PT5 using ground-based facilities is a necessary step before deciding if an ongoing mission can be extended to study a newly found NEO or perhaps a new low-cost mission is preferable.
@@ Acknowledgements.
RdlFM and CdlFM thank S. J. Aarseth for providing one of the codes used in this research and A. I. Gómez de Castro for providing access to computing facilities. JdL and JL acknowledge support from the Spanish ‘Agencia Estatal de Investigación del Ministerio de Ciencia e Innovación’ (AEI-MCINN) under the grant PID2020-120464GB-I00. This work was partially supported by the Spanish ‘Agencia Estatal de Investigación (Ministerio de Ciencia e Innovación)’ under grant PID2020-116726RB-I00 /AEI/10.13039/501100011033. Based on observations made with the Gran Telescopio Canarias (GTC), installed at the Spanish Observatorio del Roque de los Muchachos of the Instituto de Astrofísica de Canarias, on the island of La Palma. This work is partly based on data obtained with the instrument OSIRIS, built by a Consortium led by the Instituto de Astrofísica de Canarias in collaboration with the Instituto de Astronomía of the Universidad Nacional Autónoma de Mexico. OSIRIS was funded by GRANTECAN and the National Plan of Astronomy and Astrophysics of the Spanish Government. This Letter includes observations made in the Two-meter Twin Telescope (TTT) and Transient Survey Telescope (TST) in the Teide Observatory of the IAC, that Light Bridges operates in the Island of Tenerife, Canary Islands (Spain). The Observing Time Rights (DTO) used for this research were provided by PEI project SIDERA24. This Letter used flash storage and GPU computing resources as Indefeasible Computer Rights (ICRs) being commissioned at the ASTRO POC project that Light Bridges will operate in the Island of Tenerife, Canary Islands (Spain). The ICRs used for this research were provided by Light Bridges in cooperation with Hewlett Packard Enterprise (HPE) and VAST DAT. This research utilizes spectra acquired by Carle M. Pieters with the NASA RELAB facility at Brown University. In preparation of this Letter, we made use of the NASA Astrophysics Data System, the ASTRO-PH e-print server, and the MPC data server.References
@@
- Kareta et al. (2022) Kareta, T., Moskovitz, N., & Sharkey, B. 2022, AAS/DPS Meeting Abstracts, 54, 514.06
@@
- Sachdeva (2018) Sachdeva, G. S. 2018, Astropolitics, 16, 202
@@ Appendix A One Hill radius versus three Hill radii
A gravitationally bound two-body system has negative total energy. Although within the two-body problem this criterion is unambiguous, subsystems within complex -body systems may require additional constraints to ensure that a certain subsystem is indeed bound. In the particular case of planets and captured objects, Carusi & Valsecchi (1979) advocated the use of a simple criterion: the planetocentric energy of the object must be negative. Rickman & Malmort (1981) recommended adding, as second restriction, that the object completes at least one revolution around the planet. Kary & Dones (1996) recognized the critical role played by the relative distance during temporary captures by planets hosted by stars and defined temporary satellite capture by a planet using two conditions that must be met at the same time: the planetocentric energy of the object must be negative and its planetocentric distance must be under three Hill radii for the planet involved. The value of the Hill radius of Earth is roughly 0.01 au so the second condition for our planet implies flybys closer than 0.03 au. This criterion involving both geocentric energy and separation was used by Granvik et al. (2012) and Fedorets et al. (2017) to define what temporarily captured natural irregular satellites of Earth are. From there, they designated captured objects that do not complete at least one revolution around Earth as temporarily captured flybys and those completing one or more revolutions as temporarily-captured orbiters. In this work, we use the definitions in Kary & Dones (1996), Granvik et al. (2012) and Fedorets et al. (2017). For a temporary capture to take place, the intruding NEO must approach at close range (0.03 au) and low relative velocity (1 km s-1).
The Hill radius is often used to define the sphere of influence of a planet within the context of the two-body problem and the restricted three-body problem (Murray & Dermott 1999). It is sometimes argued that true temporary captures can only take place when the planetocentric energy becomes negative inside the Hill radius of the planet. We believe that the restrictive use of the one Hill radius versus the three Hill radii criterion may lead to inconsistent interpretations of true temporary orbital status of these objects. In fact, if calculations are repeated giving negligible mass to the Earth–Moon system no looping around Earth is observed; in other words, the gravitational influence of Earth in these cases is not negligible beyond one Hill radius. In addition, the strict one revolution around Earth condition cited by some is often difficult to assess because the geocentric loops traveled during temporary captures are far from closed elliptical orbits.
@@ Appendix B Mini-moons: the tally so far
So far, temporary captures by Earth, meeting the conditions in Kary & Dones (1996), Granvik et al. (2012) and Fedorets et al. (2017), have been reported for Apollo-class asteroids 2006 RH120 (Kwiatkowski et al. 2009), 2020 CD3 (Bolin et al. 2020; de la Fuente Marcos & de la Fuente Marcos 2020), 2022 NX1 (de la Fuente Marcos & de la Fuente Marcos 2022; de la Fuente Marcos et al. 2023), and 2024 PT5 (de la Fuente Marcos & de la Fuente Marcos 2024). The condition requiring negative planetocentric energy during capture is equivalent to having planetocentric eccentricity 1. Figures 5 to 8 are similar to Fig. 9 but they show the ephemerides retrieved from JPL’s Horizons instead of displaying the results of our computer simulations. They clearly show the differences between temporarily-captured orbiters (2006 RH120 and 2020 CD3) and temporarily captured flybys (2022 NX1 and 2024 PT5), but they also show that the strict application of the one Hill radius criterion may lead us to reclassify the capture episodes of 2006 RH120 and 2020 CD3 as a sequence of temporarily captured flybys, which is obviously incorrect.
@@ Appendix C Input data and uncertainties
Here, we include the barycentric Cartesian state vector of NEO 2024 PT5. This vector and its uncertainties have been used to perform the calculations discussed above and to generate the figures that display the time evolution of the critical angle, , and the geocentric energy and distance. As an example, a new value of the -component of the state vector was computed using , where is an univariate Gaussian random number, and and are the mean value and its 1 uncertainty in Table 2.
@@
| (au) | = | 9.0234018277509992.42255220 |
| (au) | = | 4.0001870467248861.26916846 |
| (au) | = | 2.6053151788957561.43228523 |
| (au/d) | = | 7.2758556060015816.00661110 |
| (au/d) | = | 1.5789959071111641.62598869 |
| (au/d) | = | 8.6555139017843991.84634694 |
Figure 9 shows the capture episode for the nominal orbit in greater detail. The minimum value of the geocentric energy coincides with a local maximum in the geocentric distance. The capture takes place when the asteroid approaches the “apogee” of its highly asymmetric geocentric trajectory and its geocentric velocity reaches a minimum.
Figure 10 is analogous to Fig. 1 but corresponds to control orbits with state vectors separated , , and from that of the nominal orbit. Results are consistent with those presented in Sect. 3.1, the orbital evolution of 2024 PT5 over the time interval (87, 61) yr is robust but outside this interval, the evolution has to be discussed in statistical terms as the evolution of arbitrarily close orbits diverges exponentially.
Appendix D TTT and TST astrometry
Astrometric observations were obtained using three telescopes located at Teide Observatory (OT, Tenerife, Canary Islands, Spain), the TTT1 and TTT2 (Two-meter Twin Telescope), and the TST (Transient Survey Telescope) as part of the key project observations during the commissioning phase. The observational circumstances are shown in Table 3. A stacked image of 2024 PT5 obtained with TTT1 is shown in Fig. 11.

### v3

@@ 3.3 Light curve
The OSIRIS instrument was also used to study the rotational properties of the object. On September 28, 2024, a series of continuous 30 s exposures was taken over one hour using the Sloan- filter and the standard CCD configuration ( binning, plate scale of 0.254 ″ pixel-1, readout time of 21 s). At the time of observation, the object had an apparent magnitude of and a total apparent angular rate of 1.2 ″ min-1. Standard bias and flat-field corrections were applied to reduce the images.
Relative aperture photometry was performed on the images using three field stars as reference stars, scaled by the mean magnitude of the target during the observed interval. Different aperture radii were tested, including both fixed and variable apertures with aperture corrections. The best result, obtained using a fixed aperture radius of about FWHM, is shown in Fig. 4.
The phased light curve is shown in Fig. 4 for a possible value of the period of 212 min (see further details in Appendix D). A variability with an amplitude of 0.3–0.4 mag was observed during the one-hour observing block, which is consistent with the uncertainty in in Table 1. We also note that the low signal-to-noise ratio (S/N) of the data do not allow an unambiguous determination of the rotation period; but our data, together with those in Bolin et al. (2025), suggest a rotation period 1 h.
@@ 4 Discussion
Our spectroscopic results on the surface composition of 2024 PT5 are not conclusive due to the lack of a near-infrared reflectance spectra, but they are suggestive of a lunar origin. Kareta et al. (2025) presented a visible to near-infrared spectrum of 2024 PT5 that, as in our case, is well matched by samples of the Moon, and is indicative of the presence of less weathered lunar-like silicates for 2024 PT5 than for (469219) 2016 HO3 Kamo’oalewa and 2022 NX1. Everything considered, the three objects may all have a common origin as lunar ejecta.
The topic of the dynamics of lunar ejecta in the neighborhood of the Earth–Moon system was first studied by Warren (1994) and Gladman et al. (1995), who arrived at the conclusion that these materials can be scattered out of their source region relatively quickly, on an effective timescale of 105 yr. This subject is now receiving renewed attention (Castro-Cisneros et al. 2023, 2025; Jiao et al. 2024) after the confirmation of Kamo’oalewa as lunar ejecta, but it is also being linked to the origin of the NEOs involved in temporary capture events (Jedicke et al. 2024). Nevertheless, the theoretical expectations drawn from the literature indicate that Kamo’oalewa, 2022 NX1, and 2024 PT5 may have all emerged out of lunar impact craters formed during the last one million years or so.
Our rotation period results are consistent with the 20 min periodic variation reported by Bolin et al. (2025). The different results obtained by Kareta et al. (2025) can be attributed to changed viewing geometry. The obtained periodic variability would be similar to that of Kamo’oalewa, which has an estimated rotation period of min (Sharkey et al. 2021). For an object with an absolute magnitude of 27.4, the possibility of a higher rotational frequency or even tumbling motion cannot be excluded. However, observations with higher time resolution and better S/N are required to confirm any such behavior.
@@ 5 Summary and conclusions
In this Letter we presented a detailed analysis of a mini-moon-like engagement experienced by Arjuna asteroid 2024 PT5 in 2024, from September 29 to November 25. The study of the short-term orbital evolution of 2024 PT5 was carried out using direct -body simulations. We derived its spectral class from a visible reflectance spectrum obtained with the OSIRIS camera spectrograph at the 10.4 m Gran Telescopio Canarias. Data from the same instrument were used to perform a preliminary assessment of its rotational state. Astrometry was obtained using the Two-meter Twin Telescope (see Appendix E). Our conclusions can be summarized as follows:
@@
We confirm that 2024 PT5 is a natural object with a visible spectrum consistent with that of lunar mare breccia, suggesting the presence of less weathered lunar-like silicates than in the case of (469219) 2016 HO3 Kamo’oalewa or 2022 NX1.
@@
We confirm that 2024 PT5 had negative geocentric orbital energy while inside a geocentric distance under three Hill radii between September 29 and November 25, 2024. We call such engagements quasi-mini-moon events.
@@
The orbital evolutions of 2022 NX1 and 2024 PT5 as Arjunas are very similar.
Both ESA and NASA now emphasize a low-cost approach to NEO missions for small body science and planetary defense, including the reuse and repurposing of already active missions (Freeman et al. 2024). The discovery and characterization of accessible objects such as 2022 NX1 and 2024 PT5 using ground-based facilities is a necessary step before deciding whether an ongoing mission can be extended to study a newly found NEO or if a new low-cost mission is perhaps preferable.
@@ Acknowledgements.
We thank the referee, M. Granvik, for a constructive, detailed, and actionable report. JdL and JL acknowledge support from the Spanish ‘Agencia Estatal de Investigación del Ministerio de Ciencia e Innovación’ (AEI-MCINN) under the grant PID2020-120464GB-I00. This work was partially supported by the Spanish ‘Agencia Estatal de Investigación (Ministerio de Ciencia e Innovación)’ under grant PID2020-116726RB-I00 /AEI/10.13039/501100011033. Based on observations made with the Gran Telescopio Canarias (GTC), installed at the Spanish Observatorio del Roque de los Muchachos of the Instituto de Astrofísica de Canarias, on the island of La Palma. This work is partly based on data obtained with the instrument OSIRIS, built by a Consortium led by the Instituto de Astrofísica de Canarias in collaboration with the Instituto de Astronomía of the Universidad Nacional Autónoma de Mexico. OSIRIS was funded by GRANTECAN and the National Plan of Astronomy and Astrophysics of the Spanish Government. This Letter includes observations made in the Two-meter Twin Telescope (TTT) and Transient Survey Telescope (TST) in the Teide Observatory of the IAC, that Light Bridges operates in the Island of Tenerife, Canary Islands (Spain). The Observing Time Rights (DTO) used for this research were provided by PEI project SIDERA24. This Letter used flash storage and GPU computing resources as Indefeasible Computer Rights (ICRs) being commissioned at the ASTRO POC project that Light Bridges will operate in the Island of Tenerife, Canary Islands (Spain). The ICRs used for this research were provided by Light Bridges in cooperation with Hewlett Packard Enterprise (HPE) and VAST DAT. This research utilizes spectra acquired by Carle M. Pieters with the NASA RELAB facility at Brown University. In preparation of this Letter, we made use of the NASA Astrophysics Data System, the ASTRO-PH e-print server, and the Minor Planet Center (MPC) data server.References
@@
- Castro-Cisneros et al. (2025) Castro-Cisneros, J. D., Malhotra, R., & Rosengren, A. J. 2025, Icarus, 429, 116379
@@
- Granvik et al. (2013) Granvik, M., Jedicke, R., Bolin, B., et al. 2013, Earth’s Temporarily-Captured Natural Satellites – The First Step towards Utilization of Asteroid Resources in Asteroids: Prospective Energy and Material Resources, ed. V. Badescu (Springer, Berlin, Heidelberg), 151
@@
- Kareta et al. (2025) Kareta, T., Fuentes-Muñoz, O., Moskovitz, N., et al. 2025, ApJ, 979, L8
@@
- Sercel et al. (2018) Sercel, J. C., Peterson, C. E., Britt, D. T., et al. 2013, Practical Applications of Asteroidal ISRU in Support of Human Exploration in Primitive Meteorites and Asteroids, ed. N. Abreu (Elsevier, Amsterdam), 477
@@ Appendix A One Hill radius versus three Hill radii
A gravitationally bound two-body system has negative total energy. Although within the two-body problem this criterion is unambiguous, subsystems within complex -body systems may require additional constraints to ensure that a certain subsystem is indeed bound. In the particular case of planets and captured objects, Carusi & Valsecchi (1979) advocated the use of a simple criterion, that the planetocentric energy of the object must be negative. Rickman & Malmort (1981) recommended adding, as a second restriction, that the object completes at least one revolution around the planet.
Kary & Dones (1996) recognized the critical role played by the relative distance during temporary captures by planets hosted by stars and defined temporary satellite capture by a planet using two conditions that must be met at the same time: The planetocentric energy of the object must be negative and its planetocentric distance must be under three Hill radii for the planet involved. The value of the Hill radius of Earth is roughly 0.01 au so the second condition for our planet implies flybys closer than 0.03 au. This criterion involving both geocentric energy and separation was used by Granvik et al. (2012) to define what temporarily captured natural irregular satellites of Earth are. From there, they designated captured objects that do not complete at least one revolution around Earth as temporarily captured flybys and those completing one or more revolutions as temporarily captured orbiters. However, Fedorets et al. (2017) made the temporary capture criterion more restrictive by decreasing the separation threshold from three Hill radii to one Hill radius. For Fedorets et al. (2017), temporarily captured flybys must have negative planetocentric orbital energy and maximum planetocentric distance under one Hill radius.
In this work, we use the definitions in Kary & Dones (1996) and Granvik et al. (2012). For a temporary capture to take place, the intruding NEO must approach at close range (0.03 au) and low relative velocity (1 km s-1). For Fedorets et al. (2017), the unusual choreography interpreted by 2024 PT5 and the Earth–Moon system is not a temporary capture.
The Hill radius is often used to define the sphere of influence of a planet within the context of the two-body problem and the restricted three-body problem (Murray & Dermott 1999). As in Fedorets et al. (2017), it is often argued that true temporary captures can only take place when the planetocentric energy becomes negative inside the Hill radius of the planet. We believe that the restrictive use of the one Hill radius versus the three Hill radii criterion may lead to inconsistent interpretations of the actual temporary orbital status of these objects. In fact, if calculations are repeated giving negligible mass to the Earth–Moon system, no looping around Earth is observed; in other words, the gravitational influence of Earth in these cases is not negligible beyond one Hill radius. In addition, the strict one revolution around Earth condition cited by some is often difficult to assess because the geocentric loops traveled during temporary captures are far from closed elliptical orbits.
@@ Appendix B Mini-moons: The tally so far
So far, temporary captures by Earth, meeting the conditions in Kary & Dones (1996) and Granvik et al. (2012), have been reported for Apollo-class asteroids 2006 RH120 (Kwiatkowski et al. 2009), 2020 CD3 (Bolin et al. 2020; de la Fuente Marcos & de la Fuente Marcos 2020), 2022 NX1 (de la Fuente Marcos & de la Fuente Marcos 2022; de la Fuente Marcos et al. 2023), and 2024 PT5 (de la Fuente Marcos & de la Fuente Marcos 2024; Bolin et al. 2025). However, considering the more restrictive criterion featured in Fedorets et al. (2017), only 2006 RH120, 2020 CD3, and 2022 NX1 experienced temporary captures by Earth.
The condition requiring negative planetocentric energy during capture is equivalent to having planetocentric eccentricity 1. Figures 5 to 8 are similar to Fig. 9, but they show the ephemerides retrieved from JPL’s Horizons instead of displaying the results of our computer simulations. They clearly show the differences between temporarily captured orbiters (2006 RH120 and 2020 CD3) and temporarily captured flybys (2022 NX1 and 2024 PT5), but they also show that the strict application of the one Hill radius criterion may lead us to reclassify the capture episodes of 2006 RH120 and 2020 CD3 as a sequence of temporarily captured flybys, which is obviously incorrect.
@@ Appendix C Input data and uncertainties
Here we include the barycentric Cartesian state vector of NEO 2024 PT5. This vector and its uncertainties were used to perform the calculations discussed above and to generate the figures that display the time evolution of the critical angle, , and the geocentric energy and distance. As an example, a new value of the -component of the state vector was computed using , where is an univariate Gaussian random number, and and are the mean value and its 1 uncertainty in Table 2.
@@
| (au) | = | 9.0234022187771964.92897003 |
| (au) | = | 4.0001870598088878.32396963 |
| (au) | = | 2.6053077272983945.84709025 |
| (au/d) | = | 7.2758546483671929.95845747 |
| (au/d) | = | 1.5789959078314936.93400853 |
| (au/d) | = | 8.6554184539657677.24496957 |
Figure 9 shows the capture episode for the nominal orbit in greater detail. The minimum value of the geocentric energy coincides with a local maximum in the geocentric distance. The capture takes place when the asteroid approaches the apogee of its highly asymmetric geocentric trajectory and its geocentric velocity reaches a minimum.
Figure 10 is analogous to Fig. 1, but corresponds to control orbits with state vectors separated by , , and from that of the nominal orbit. These results are consistent with those presented in Sect. 3.1, the orbital evolution of 2024 PT5 over the time interval (87, 60) yr is robust, but outside this interval the dynamics have to be discussed in statistical terms as the evolution of arbitrarily close orbits diverges exponentially. Figure 10, top panel, shows that 2024 PT5 might have experienced 1:1 resonant behavior of the horseshoe type in the past, and perhaps it will again in the future, but the evolution is strongly affected by orbital chaos driven by close encounters with the Earth–Moon system.
Appendix D Time series analysis
We analyzed the light curve using the Phase Dispersion Minimization method (PDM; Stellingwerf 1978) to find its rotation period. The PDM method minimizes the sum of the squares of the differences in the ordinate from one data point to the next. The chosen period is the one resulting in the smallest sum; in other words, the one producing the least possible scatter about the derived light curve. The dispersion as a function of the test period is shown in Fig. 11. We obtained a possible variation with a period =212 min.
Appendix E TTT and TST astrometry
Astrometric observations were obtained using three telescopes located at Teide Observatory (OT, Tenerife, Canary Islands, Spain), the TTT1 and TTT2 (Two-meter Twin Telescope), and the TST (Transient Survey Telescope) as part of the key project observations during the commissioning phase. The observational circumstances are shown in Table 3. Additional astrometry from these telescopes was submitted to the Minor Planet Center (MPC).111111https://minorplanetcenter.net/db_search/show_object?utf8=%E2%9C%93&object_id=2024+PT5 A stacked image of 2024 PT5 obtained with TTT1 is shown in Fig. 12.